OpenStudy (mmswk):

PLEASE HELP ME! I HAVE NEVER SEEN THIS BEFORE AND I NEED TO GET IT RIGHT! I GIVE MEDALS!

OpenStudy (mmswk):

\[10/3/1/6 - (2.8 + 1.2)\] The 10/3 is over a / and then 1/6 on the bottom

OpenStudy (reemii):

\[ \frac{\frac{10}{3}}{\frac16} - (2.8 + 1.2) \] Use the formula: \( \frac{\frac{a}b}{\frac{c}d} = \frac{ad}{bc} \) (It all follows from the fact that \(\frac{X}{Y} = X\frac1Y\). )

OpenStudy (mmswk):

??????

OpenStudy (mmswk):

is it 26?

OpenStudy (fluttershyk):

@rebeccaxhawaii can you help him/her

OpenStudy (reemii):

Yes. \[ \frac{\frac{a}b}{\frac{c}d} = \frac{ad}{bc} \] So, \[ \frac{\frac{10}3}{\frac{1}6} = \frac{10}{3}\frac{6}{1} = \frac{60}{3}\] Subtract (2.8+1.2) from \(60/3\).

OpenStudy (reemii):

* not "yes" ..

OpenStudy (mmswk):

no or yes?

OpenStudy (reemii):

It is not 26. What is 60/3? What 2.8+1.2?

OpenStudy (mmswk):

wait 20 right?

OpenStudy (reemii):

yes (60/3 = 20)

OpenStudy (mmswk):

Thank you so much!!!

OpenStudy (reemii):

don't forget to remove 4 from 20. and you have the answer.